OpenStudy (anonymous):

reduce -56/-70 to lowest terms a)8/10 b)-4/5 c)-2/3 d)4/5

OpenStudy (whpalmer4):

are there any numbers which you can divide both -50 and -70 by and have it come out without a remainder?

OpenStudy (whpalmer4):

might help to think of it as factors: -50 = -1 * 2*5*5 -70 = -1*2*5*7 \[\frac{-1*2*5*5}{-1*2*5*7} = \frac{\cancel{-1*2*5}*5}{\cancel{-1*2*5}*7}=\frac{5}{7}\]

OpenStudy (whpalmer4):

you've got the wrong answer choices up there...

OpenStudy (anonymous):

sorry fixed the question

OpenStudy (whpalmer4):

Okay. same approach applies.

OpenStudy (whpalmer4):

-56 = -1 * 2* 2*2*7 -70 = -1*2*5*7 \[\frac{-56}{-70}= \frac{-1*2*2*2*7}{-1*2*5*7} = \frac{2*2}{5} = \]

OpenStudy (anonymous):

thanks

OpenStudy (whpalmer4):

You could also do this by successively peeling off factors: clearly both have -1 as a factor, so -56/-70 = 56/70. now both are even, so we can divide both by 2: 56/70 = 28/35 both of those are divisible by 7, so 4/5 is the final result

OpenStudy (anonymous):

thanks, you were right! :)